OpenStudy (anonymous):

Lol I might jsut be stupid but idk how to do this xD Find the value of the cot 10° using your calculator. IDK HOW D; It's my last question I don't have a cot button on my calc either if I'm supposed to and and

OpenStudy (anonymous):

do tan 10 then put 1 ÷ Ans

OpenStudy (anonymous):

Firstly find tan10..

OpenStudy (anonymous):

cot=1/tg

OpenStudy (anonymous):

Or you could do 1 ÷ (tan 10)

OpenStudy (anonymous):

oh. I am stupid. My bad. Thanks guys :D

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

\[\cot \theta = \frac{1}{\tan \theta} = \frac{\cos \theta}{\sin \theta}\]

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

going hard @agentx5

OpenStudy (anonymous):

:O AGENT D: -___-

OpenStudy (anonymous):

I can make it more hard: \[\cot \theta = \frac{1}{\tan \theta} = \frac{\sin(2 \theta)}{2\sin^2 \theta}\]

OpenStudy (anonymous):

oh wait xD Is that tan^-1 or tan (10)

OpenStudy (anonymous):

lol "more hard"

OpenStudy (anonymous):

tan10..

OpenStudy (anonymous):

YES @waterineyes

OpenStudy (anonymous):

Double angles rule!

OpenStudy (anonymous):

0.176?

OpenStudy (anonymous):

@rebeccaskell94 i think now you are getting more confused..

OpenStudy (anonymous):

no no no not the same thing dear Rebecca! :-D \[\tan^{-1} (\frac{opp}{adj}) = \arctan (\frac{opp}{adj}) = \theta\]

OpenStudy (anonymous):

lol Agent your genius talents are commendable

OpenStudy (anonymous):

Now use: 1/0.176

OpenStudy (anonymous):

Inverse tangent is cotangent, anti-tangent is arctangent. The latter returns and angle value in degrees or radians :-)

OpenStudy (anonymous):

@rebeccaskell94 You forgot to do the rciprocal. Now. 1/0.176326981...

OpenStudy (anonymous):

5.671? o.o

OpenStudy (anonymous):

THat's what I get :)

OpenStudy (anonymous):

oh dang.

OpenStudy (anonymous):

May be.. Ha ha ha..

OpenStudy (anonymous):

3:)

OpenStudy (anonymous):

Can anyone find this without calculator??

OpenStudy (anonymous):

o_O No idea...

OpenStudy (anonymous):

Is Wolfram considered a calculator? XD

OpenStudy (anonymous):

Can anybody think about finding the above value without using calculator?? Question changed to this one now..

OpenStudy (anonymous):

I can... I think...